OpenStudy (anonymous):

Giving Medals ((&)) A Fan Solve -4(x + 10) - 6 = -3(x - 2). A)-40 B)-46 C)-52 D)52

OpenStudy (anonymous):

Are you familiar with the distribution principal?

OpenStudy (anonymous):

yes (-4x - 40) - 6 = (-3x +6)

OpenStudy (anonymous):

Great so your equation would look like this now -4x-40-6=-3x+6. Correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now get everything with an x on one side of the equation and all other values on the other side of the equation. Understand?

OpenStudy (anonymous):

So you would subtract -4x from 3x which gives you -7x ?

OpenStudy (anonymous):

No. Remember what ever you do to one side of the equation you must do to the other. What if you add 4x to both sides? What would your equation look like?

OpenStudy (anonymous):

would it be ... x-40-6= x + 6

OpenStudy (anonymous):

4x+-4x=0 correct so it wouldn't be x-40-6 it would be -40-6=x+6 right?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Now you need to get that +6 to the left side of the equation and you will have the value for x

OpenStudy (anonymous):

so would it be ... -40=x + 6

OpenStudy (anonymous):

No..remember what ever you do to one side you have to do to the other side. -40-6=x+6. So if I want to make the +6=0 what do I have to add to it?

OpenStudy (anonymous):

add -6 + 6

OpenStudy (anonymous):

And yu have to do the same to the otherside so we add -6 to both sides and end up with?

OpenStudy (anonymous):

do you add -6 to -40 ?

OpenStudy (anonymous):

Your equation looks like this -40-6=x+6. If I add -6 to both sides I get -40-6-6 = x+6-6. So what do you end up when you simplify>

OpenStudy (anonymous):

Ohh is the answer -52 ?

OpenStudy (anonymous):

Very good. You seem to have a grasp of what needs to be done. Just always remember what ever you do to one side of an equation you MUST do the same to the other side of the equation

OpenStudy (anonymous):

Thanks so much !